# CREStE: Scalable Mapless Navigation with Internet Scale Priors and Counterfactual Guidance

### v1

@@ Abstract
We address the long-horizon mapless navigation problem: enabling robots to traverse novel environments without relying on high-definition maps or precise waypoints that specify exactly where to navigate. Achieving this requires overcoming two major challenges — learning robust and generalizable perceptual representations of the environment without pre-enumerating all possible navigation factors and forms of perceptual aliasing, and utilizing these learned representations to plan human-aligned navigation paths. Existing solutions struggle to generalize due to their reliance on: (a) hand-curated object lists, which overlook new, unforeseen factors, (b) end-to-end learning of navigation-relevant features, which is constrained by the limited availability of real robot data, (c) large sets of expert demonstrations, which provide insufficient guidance on the most critical perceptual cues, or (d) handcrafted reward functions for learning, which are difficult to design and adapt for new scenarios. To overcome these limitations, we propose CREStE, the first method that learns representations and rewards capable of addressing the full mapless navigation problem without relying on large-scale robot datasets or manually curated features. CREStE leverages visual foundation models trained on internet-scale data to learn continuous bird’s-eye-view representations capturing elevation, semantics, and instance-level features. To utilize learned representations for planning, we propose a counterfactual-based loss and active learning procedure to focus on the perceptual cues that matter most by querying humans for counterfactual trajectory annotations in challenging scenes. We evaluate CREStE in kilometer-scale navigation tasks across six distinct urban environments. CREStE significantly outperforms all existing state-of-the-art approaches with 70% fewer human interventions per mission, including a 2-kilometer mission in an unseen environment with just 1 intervention; showcasing its robustness and effectiveness for long-horizon mapless navigation. Videos and additional materials can be found on the project page: https://amrl.cs.utexas.edu/creste.
Figure 1: CREStE predicts bird’s eye view (BEV) feature and reward maps for mapless navigation in urban environments. Our RGB-D and BEV backbone work together to predict a structured BEV feature map consisting of separate feature maps for static/dynamic entities and elevation, which our reward function uses to predict expert-aligned rewards. To learn this feature map, CREStE proposes a novel method of distilling internet scale priors from visual foundation models. We also introduce an active learning framework and training objective that improves reward alignment using counterfactual demonstrations. We integrate CREStE in a modular navigation system that uses coarse GPS guidance and reward maps to reach navigation goals safely.
@@ I Introduction
Mapless navigation is the task of reaching user-specified goals without high-definition (HD) maps and precise navigation waypoints. Instead, mapless navigation approaches rely on egocentric sensor observations (e.g. RGB images, point clouds, GPS), coarse waypoints from public routing services, and satellite imagery to plan safe paths aligned with human preferences. While this task is relevant to many settings, our work focuses on mapless navigation for challenging urban scenes like downtowns, parks, and residential neighborhoods.
Mapless navigation is desirable as it enables rapid deployment of robots to novel environments without laborious map construction and maintenance. Furthermore, solutions to mapless navigation broadly improve robustness to real-world factors that affect both map-centric and mapless methods, such as unexpected environment changes and traffic patterns for dynamic entities.
Geometric-only methods [3, 46] are one approach to mapless navigation that considers geometric factors like static obstacles. While understanding geometric factors is important, they are insufficient alone for explaining navigation preferences in diverse urban scenes that require jointly reasoning about terrain preferences (grass vs. concrete), semantic cues (crosswalks and crossing signs), and uneven terrain (sharp curb dropoffs). Furthermore, methods need to reason about the factors above without pre-enumerating the list of semantic classes and entities a priori to scale gracefully in diverse environments. Lastly, planning modules that reason about these factors must identify the most salient features and understand how they influence which paths are most likely to be aligned with human preferences.
@@
Mapless navigation approaches that consider factors beyond geometry broadly consist of single-factor perception, hand-curated multi-factor perception, end-to-end learning, and zero-shot pre-trained large language model (LLM)/visual language model (VLM) transfer. Single-factor [23, 47, 13] and multi-factor perception [24, 37] approaches encode either single or multiple factors (e.g. terrain, elevation, semantics) to a learned representation and use analytical methods like model predictive control [35] for path planning. Single-factor methods generalize poorly to complex settings, while multi-factor methods, though more flexible, rely on a hand-curated list of semantic classes and terrains, limiting generalization to unseen classes. End-to-end learning methods [38, 40, 39] jointly learn the representation and policy from expert demonstrations, but are prone to overfitting without access to large-scale robot datasets. While recent works [48, 25] show that pre-trained factors in LLMs and VLMs can be leveraged for navigation, we empirically demonstrate that they are not well-attuned to urban navigation, leading to poor zero-shot transfer performance in complex scenes.
We address the aforementioned limitations with CREStE, Counterfactuals for Reward Enhancement with Structured Embeddings, the first method that learns generalizable perceptual representations for the full mapless navigation problem in urban environments. Our key insights lie in how we leverage visual foundation models and counterfactual demonstrations to learn structured representations and policies for mapless navigation. First, CREStE learns a perceptual encoder by synergizing navigation priors from multiple visual foundation models (VFMs) trained on internet-scale data, yielding semantic, geometric, and entity aware representations that are robust to perceptual aliasing and geometrically grounded in the robot’s navigation horizon. We find it is essential to unify priors from multiple VFMs as each model is best suited for different factors: SegmentAnything [33, 20] for distinguishing between entities and CLIP [32]/Dino [5, 28] for inferring visual-language and semantic features. Second, CREStE learns a reward function for navigation using our principled active learning framework and IRL objective that queries humans for counterfactuals to align learned policies with human preferences. Counterfactuals (i.e. alternate trajectories that do not align with preferences) enable reward functions to reason about the most salient perceptual features and learn complex navigation behaviors under a unified objective, removing the need for large-scale robot datasets or careful reward design.
We summarize these contributions below:
@@
  • Counterfactually Aligned Rewards. An active learning framework and principled counterfactual IRL formulation for reward alignment using counterfactual and expert demonstrations.
We demonstrate our approach’s effectiveness through real-world kilometer-scale navigation experiments, such as the one shown in Fig. 5. In total, we evaluate our approach in 6 distinct seen and unseen urban environments, outperforming existing state-of-the-art imitation learning, inverse reinforcement learning, and heuristic-based methods on the task of mapless navigation.
@@ II Related Work
In this section, we position our work within the broader context of methods that perform mapless navigation by learning representations and policies for safe local path planning. Existing solutions broadly consist of hybrid approaches that combine single/multi-factor learned representations with analytical path planning, end-to-end methods that learn representations and policies jointly from data, and VLM/LLM-based methods that leverage large foundation models zero-shot for navigation. Underlying these approaches are two key challenges: learning robust, generalizable perceptual representations that encode a sufficient set of navigation factors, and identifying and reasoning about the important factors to plan safe paths.
@@
Single [16, 23] or multi-factor [10, 24, 37] approaches explicitly learn high-level representations that consider factors such as terrain, semantic, or geometric understanding. Using manually curated cost functions to balance these factors, these methods map representations to scalar costmaps for analytical planning and control methods like model predictive control (MPC) [35]. While these methods can jointly reason about many navigation factors, they assume the full set of semantic classes and terrains are static and known apriori, and consequently generalize poorly to unexpected semantic classes, terrains, or other factors not seen during training. Furthermore, they require expert tuning of complex multi-objective cost functions to jointly consider multiple factors.
@@
End-to-end methods learn a representation and policy together using either behavior cloning, RL, or IRL. Behavior cloning methods [17, 40] implicitly learn representations with navigation cues and how to reason about them by mimicking the expert policy. These approaches scale effectively given large-scale datasets with expert demonstrations but are prone to overfitting otherwise. RL [12] and IRL [50] methods also mimic the expert policy, but additionally leverage handcrafted reward functions (e.g. minimize vibrations [14] and likelihood of interventions [15]) to guide representation and policy learning. This improves learning efficiency at the cost of careful reward tuning to learn fine-grained behaviors that are difficult to infer from expert demonstrations alone (e.g. avoid collisions with walls but not tall grass). Reward functions also offer more interpretability, particularly during failures where reward functions make it possible to identify the root cause.
@@
An emerging class of methods [25, 48] leverage pre-trained factors present in VLMs and LLMs, large foundation models pre-trained on internet-scale data, zero-shot for navigation. These foundation models leverage geographic hints in the form of satellite imagery and topological maps, along with image observations and text instructions to predict safe local waypoints for analytical planning and control methods. While promising, it is not well understood how to best guide VLMs/LLMs to reason about the most important navigation factors for long-horizon navigation tasks.
CREStE is a hybrid approach that learns perceptual representations by distilling navigation factors from VFMs.pre-trained on internet-scale data and leverages IRL to efficiently learn expert-aligned behavior. This stands in contrast to prior work that either learn the representation from expert demonstrations or bespoke datasets, and approaches that extract navigation factors from VLMs/LLMs in a zero-shot manner. This provides a number of benefits when scaling to diverse urban settings: 1) Improving robustness to perceptual aliasing due to lighting, weather, and viewpoint variations; 2) Promoting generalizability to unseen semantic entities; 3) Encoding semantic and entity aware priors without dense human labels. Additionally, CREStE replaces complex multi-objective reward functions common for IRL and RL methods with a unified counterfactual-based learning objective. Counterfactual demonstrations provide a theoretically principled way to communicate complex navigation operator preferences that would be otherwise difficult to specify from expert demonstrations alone.
@@ III-A Path Planning for Mapless Navigation
$\Gamma_{S}=\text{arg}_{\Gamma}\text{min}||x_{t+S}-G||+\lambda_{t}\mathcal{J}(\Gamma),$ (1)
where $||x_{t+S}-G||$ is the distance between the final state $x_{t+S}$ and $G$, and $\mathcal{J}(\Gamma)$ is a cost function for path planning scaled by a relative weight $\lambda_{t}$. In mapless urban environments, the robot pose $x_{t}$ and goal $G$ may be highly noisy, causing $\mathcal{J}(\Gamma)$’s importance to vary dynamically across time. In this work, we do not address this problem and assume $\lambda_{t}$ to be static.
@@ III-B General Preference-Aligned Cost Functions
To properly define our cost function $\mathcal{J}(\Gamma)$, we first need to introduce the notion of an observation function $\Theta:\mathcal{O}\rightarrow\mathcal{T}$ that maps observations $o_{t}$ to a joint distribution $\mathcal{T}$ that encapsulates a set of relevant factors for navigation in the world. In general, the sufficient set of factors is unknown but can be approximated for each environment. Let $T:\mathcal{X}\rightarrow\mathcal{T}$ be a function that maps a robot pose $x\in\mathcal{X}$ to a feature $\tau\in\mathcal{T}$, where $\tau$ captures the relevant set of factors necessary to reason about $\mathcal{J}(\Gamma)$. The relationship and set of relevant factors may consist of geometric, semantic, and social costs as follows:
@@
where $\sigma$ is a nonlinear function that combines individual cost terms.
@@ IV Approach
Figure 2: Training procedure and model architecture for the CREStE perceptual encoder $\Theta$. Our RGB-D backbone extracts semantic features and performs depth completion from a single RGB and sparse depth image pair. Next, we lift and splat the image features $z_{\text{rgbd}}$ to an unstructured BEV feature map before predicting continuous static panoptic, dynamic panoptic, and elevation feature maps. Finally, we stack the predicted BEV features to construct a structured BEV representation for our learned reward function. We supervise $\Theta$ using semantic feature maps, completed depth, and BEV map labels generated by our SegmentAnythingv2 [20] and Dinov2 [28]-powered distillation label generator. We define the dimensions for important feature maps in the Feature Dimension Legend on the bottom right.
We now present CREStE, an end-to-end learning-based system that predicts general-purpose, egocentric reward maps for local path planning in complex, urban environments. Unlike prior works [40, 49] that learn solely from expert demonstrations, our method leverages large visual foundation models (VFMs) to learn robust perceptual representations rich with navigation priors and counterfactual demonstrations to align reward maps more closely with operator preferences.
CREStE is a modular approach with two key components: 1) A perceptual encoder $\Theta(o_{\text{rgb, t}}$, $o_{\text{depth, t}})$ that takes the robot’s current RGB and sparse depth observation and predicts a completed depth image $y_{\text{depth, t}}$ and structured BEV feature map $y_{\text{bev, t}}$; 2) A reward function $r_{\phi}(y_{\text{bev, t}})$ that takes $y_{\text{bev, t}}$ and outputs a BEV scalar reward map $y_{\text{reward, t}}$. $\Theta$ captures key semantic, geometric, and entity factors crucial for urban navigation and $r_{\phi}$ jointly reasons about these factors to produce local BEV reward maps for local path planning. Both components are learned from VFMs, expert demonstrations, and offline counterfactual annotations, without relying on dense human labels, exhaustive semantic categories, or large-scale expert datasets.
In the remainder of this section, we describe the following: 1) Sec. IV-A - The CREStE model architecture, highlighting key innovations for learning the perceptual encoder $\Theta$ in Sec. IV-A1 and reward function $r_{\phi}$ in Sec. IV-A2. 2) Sec. IV-B - The CREStE training procedure, where Sec. IV-B1 presents our learning and VFM distillation procedure for $\Theta$ and Sec. IV-B2 presents our active reward learning framework for learning $r_{\phi}$ from expert and counterfactual demonstrations.
@@ IV-A CREStE Model Architecture
CREStE implements $\Theta$ and $r_{\phi}$ using four major components. Fig. 2 illustrates the first three components that constitute $\Theta$: 1) RGB-D encoder ($f_{\text{rgbd}}$) encodes important semantic cues for navigation (e.g. crosswalk markings, objects, terrains) and geometric information (dense depth), 2) Lift-splat module ($f_{\text{splat}}$) lifts and grounds image features into a local BEV map, and 3) BEV Inpainting Backbone ($f_{\text{bev}}$) completes the scene, predicting a structured BEV feature map $y_{\text{bev}}$ with continuous entity-aware semantic and elevation layers. The fourth component is our reward function $r_{\phi}(y_{\text{bev}})$, which consumes the BEV representation $y_{\text{bev}}$ and predicts a scalar reward map $y_{\text{reward}}$ for navigation. We define specific details regarding the dimensions of each feature in Fig. 2 and present our architecture and innovations for $\Theta$ in Sec. IV-A1 and $r_{\phi}$ in Sec. IV-A2.
@@ IV-A1 Perceptual Encoder Model Architecture
Our 25.5M-parameter perceptual encoder draws inspiration from the TerrainNet [24] architecture while introducing two key model innovations that improve generalizability to diverse urban scenes. These innovations are our semantic decoder head $f_{\text{semantic}}$ and BEV map decoders ($f_{\text{bev,static}}$, $f_{\text{bev,dynamic}}$) that predict continuous panoptic features for static and dynamic entities. We borrow the term “panoptic” [19] from the computer vision community, where it refers to the task of understanding both the semantics and instances in the scene. We present an overview of these components next.
RGB-D Encoder. $f_{\text{rgbd}}(o_{\text{rgb}},o_{\text{depth}})$ processes the RGB and sparse depth image using an EfficientNet-B0 [45] encoder and outputs a latent feature map $z_{\text{rgbd}}$. We pass $z_{\text{rgbd}}$ to two decoder heads: 1) a depth completion head $f_{\text{depth}}(z_{\text{rgbd}})$ that outputs a completed depth map $y_{\text{depth}}$ and 2) a semantic decoder head $f_{\text{semantic}}(z_{\text{rgbd}})$ that regresses a semantic feature map $y_{\text{semantic}}$.
Lift-Splat Module. Adopting TerrainNet’s soft-quantization approach, $f_{\text{splat}}(z_{\text{rgbd}},y_{\text{depth}})$ lifts the output of $f_{\text{rgbd}}$ to 3D space using $y_{\text{depth}}$ and “splats” them onto a BEV grid to produce an unstructured feature map $z_{\text{bev, splat}}$. We learn weights to bilinearly interpolate each image feature to its corresponding cell and 4 neighboring map cells, which simultaneously grounds features to the robot’s local planning horizon and corrects depth prediction inconsistencies.
BEV Inpainting Backbone. We implement the final component $f_{\text{bev}}(z_{\text{bev, splat}})$ using a U-Net [36] architecture with a shared BEV map encoder $f_{\text{bev, enc}}$ and three decoder heads, $f_{\text{bev, static}}(z_{\text{bev, enc}})$, $f_{\text{bev, dynamic}}(z_{\text{bev, enc}})$, and $f_{\text{bev, elev}}(z_{\text{bev, enc}})$, where $z_{\text{bev, enc}}$ is the latent feature map produced by $f_{\text{bev, enc}}$. Our three decoder heads predict static panoptic feature maps $y_{\text{bev, static}}$, dynamic panoptic feature maps $y_{\text{bev, dynamic}}$, and elevation maps $y_{\text{bev, elev}}$ respectively. We stack these BEV feature maps along the channel dimension to construct a structured BEV feature map $y_{\text{bev}}$ for downstream reward learning.
Design Rationale and Key Innovations. Our architectural innovations aim to address three challenges to perceptual representation learning for mapless urban navigation: 1) Robustness to perceptual aliasing, where representations are lighting, viewpoint, and weather invariant; 2) Generalization beyond the training data to unseen semantic classes and terrains; 3) Panoptic understanding without exhaustively enumerating all possible semantic classes and terrains a priori.
We address challenges 1 and 2 by using $f_{\text{semantic}}$ to learn representations that regress to the features produced by large-scale Visual Foundation Models (VFMs) such as Dinov2 [28]. By matching Dinov2’s learned representations given the same RGB observations, our model inherits robust semantic and geometric priors. Furthermore, these priors capture high-level semantic cues without restricting the model to a fixed set of semantic classes. As a result, $\Theta$ is expected to generalize beyond the classes and terrain types present in the robot’s training data.
However, Dinov2 alone lacks the entity-awareness needed to accurately project and inpaint features in heavily occluded urban scenes, particularly when working with noisy and sparse depth inputs. $f_{\text{bev, static}}$ and $f_{\text{bev, dynamic}}$ remedy this by learning a map representation aligned with BEV instance labels derived from SegmentAnythingv2 (SAM2) [20]. This ensures our model can identify and localize individual entities—both static and dynamic—within a local BEV horizon, thus bolstering its capacity to handle complex occlusions and ambiguous depth cues. Working together, our decoder heads synergistically balance learning a perceptual representation that is simultaneously semantic, entity-aware, and geometrically discriminative, all while maintaining strong robustness to perceptual aliasing.

### v2

@@ Abstract
We introduce CREStE, a scalable learning-based mapless navigation framework to address the open-world generalization and robustness challenges of outdoor urban navigation. Key to achieving this is learning perceptual representations that generalize to open-set factors (e.g. novel semantic classes, terrains, dynamic entities) and inferring expert-aligned navigation costs from limited demonstrations. CREStE addresses both these issues, introducing 1) a visual foundation model (VFM) distillation objective for learning open-set structured bird’s-eye-view perceptual representations, and 2) counterfactual inverse reinforcement learning (IRL), a novel active learning formulation that uses counterfactual trajectory demonstrations to reason about the most important cues when inferring navigation costs. We evaluate CREStE on the task of kilometer-scale mapless navigation in a variety of city, offroad, and residential environments and find that it outperforms all state-of-the-art approaches with 70% fewer human interventions, including a 2-kilometer mission in an unseen environment with just 1 intervention; showcasing its robustness and effectiveness for long-horizon mapless navigation. Videos and additional materials can be found on the project page: https://amrl.cs.utexas.edu/creste.
Figure 1: CREStE learns a perceptual encoder1,2 and reward function3 to predict structured bird’s eye view (BEV) feature and reward maps for navigation. Our model inherits generalization and robustness from visual foundation models and learns expert-aligned rewards using our counterfactual-guided learning framework. We integrate CREStE in a modular navigation system that uses coarse GPS guidance and rewards to reach navigation goals safely.
@@ I Introduction
Mapless navigation is the task of reaching user-specified goals without high-definition (HD) maps and precise navigation waypoints. Mapless approaches plan routes using egocentric sensor observations (e.g. RGB images, point clouds, GPS), coarse waypoints from public routing services, and satellite imagery. These solutions demonstrate promise as a scalable alternative to conventional map-centric approaches: enabling generalization to unforeseen factors (e.g. curb ramps and foliage) and dynamic entities (e.g. pedestrians and strollers) while reducing map maintenance overhead and reliance on pre-defined routes.
Traditional geometric-only mapless solutions [3, 43] exhibit robust generalization when it is only necessary to consider costs for geometric factors like static obstacles. However, open-world navigation demands perception systems that perceive an open set of factors unknown apriori, ranging from terrain preferences (grass vs. concrete) to semantic cues (crosswalks and crossing signs). Furthermore, it requires the ability to identify the most salient features in the scene, and how they influence the choice of paths.
@@
Learning-based approaches are a scalable alternative that considers factors beyond geometry, but must overcome data scarcity, robustness, and generalization challenges to achieve similar reliability. These approaches broadly consist of single-factor perception, hand-curated multi-factor perception, end-to-end learning, and zero-shot pre-trained large language model (LLM)/visual language model (VLM) transfer. While single-factor [20, 44, 43] and multi-factor [21, 33] methods learn representations that consider relevant navigation factors (e.g. geometry, terrain, semantics), they rely on a hand-curated list of semantic classes and terrains that limit generalization to unseen classes. End-to-end methods [34, 35, 36] alleviate this by jointly learning the representation and policy from expert demonstrations, but are prone to overfitting without large-scale robot datasets. While recent works [45, 22] demonstrate that pre-trained LLMs and VLMs can reason about expert-aligned behavior without large-scale robot datasets, we empirically demonstrate that they are poorly attuned to urban navigation, leading to brittle zero-shot transfer in complex scenes.
We address the aforementioned limitations with CREStE, Counterfactuals for Reward Enhancement with Structured Embeddings, a novel approach that learns open-set representations and navigation costs/rewards for mapless urban navigation. To learn open-set representations, CREStE combines prior work on bird’s eye view (BEV) representation learning [21] and visual foundation models (VFMs) [25, 30] to learn BEV map representations with inherited real-world robustness and open-set semantic knowledge. We unify priors from multiple foundation models by distilling image features from Dinov2 [25] and refining these features in BEV using SAM2 [30] instance labels. CREStE infers expert-aligned navigation costs without large-scale demonstrations by leveraging counterfactuals to guide learning, where counterfactuals hold all other variables constant except for the path taken from the start to the end goal. Unlike conventional preference learning, our proposed counterfactual inverse reinforcement learning (IRL) objective explicitly minimizes rewards along paths that exhibit undesirable behavior (e.g. veering off crosswalks, driving off curb ramps, etc.), enabling operators to correct robot behavior by providing offline counterfactual feedback. We summarize the main contributions of our work as follows:
@@
  • Counterfactually Aligned Rewards. An active learning framework and counterfactual IRL formulation for reward alignment using counterfactual and expert demonstrations.
We demonstrate our approach’s effectiveness through real-world kilometer-scale navigation experiments in urban environments with off-road terrain, elevated walkways, sidewalks, intersections, and other challenging conditions. In total, we evaluate in 6 distinct seen and unseen geographic areas and significantly outperform existing state-of-the-art imitation learning, inverse reinforcement learning, and heuristic-based methods on the task of mapless navigation.
@@ II Related Work
In this section, we position our work within the broader context of methods that perform mapless navigation by learning representations and policies for safe local path planning. Underlying these methods are two key challenges: learning robust perceptual representations that encode an open set of navigation factors, and learning planning modules that can identify and reason about the most important factors to plan safe paths.
@@
Single [14, 20, 39] and multi-factor [9, 21, 33] approaches learn perceptual representations grounded in the planning horizon with elevation, terrain, and semantic factors, and rely on expert tuning to balance the relative costs of each factor. While this enables joint reasoning about multiple factors, it assumes the full set of semantic classes and terrains are known apriori, generalizing poorly to unexpected semantic classes, terrains, or other factors not seen during training. Furthermore, they require expert tuning of complex multi-objective cost functions for each environment.
@@
End-to-end methods jointly learn a representation and policy using either behavior cloning (BC), RL, or IRL. BC methods [15, 35] learn task-specific representations for reasoning about how to mimic the expert policy. These approaches scale effectively given large-scale datasets with expert demonstrations, but are prone to overfitting otherwise. RL [11] and IRL [47] methods also mimic the expert policy, but design handcrafted reward functions (e.g. minimize vibrations [12] and likelihood of interventions [13]) or preference rankings [40, 7] to guide representation and policy learning. These methods efficiently learn expert-aligned behaviors at the cost of careful reward tuning or preference enumeration.
@@
An emerging class of methods [22, 45] leverages pre-trained factors present in VLMs and LLMs, large foundation models pre-trained on internet-scale data, zero-shot for navigation. These foundation models leverage geographic hints in the form of satellite imagery and topological maps, along with image observations and text instructions to predict safe local waypoints for analytical planning and control methods. While promising, it is not well understood how to best steer VLMs/LLMs to reason about the most important navigation factors for navigation tasks.
The aforementioned works either rely on large-scale robot datasets for open-set generalization or complex reward design to learn expert-aligned behavior. However, constructing large robot datasets and expressive multi-term objectives is prohibitively difficult for open-world settings. In contrast, CREStE learns open-set representations by directly distilling knowledge from VFMs, and expert-aligned reward functions by using counterfactuals as a unified language for conveying navigation constraints. Unlike counterfactual-based methods like VRL-PAP [39], our counterfactuals specify arbitrary navigation constraints beyond terrains and can be readily obtained offline, enabling a data flywheel for continuous improvement with additional deployments. Importantly, our reward learning approach is distinct from preference-learning methods, as counterfactuals are a more specific form of feedback than preferences, which can be between any two trajectories.
@@ III-A Path Planning for Mapless Navigation
$\Gamma_{S}=\text{arg}_{\Gamma}\text{min}||x_{t+S}-G||+\lambda\mathcal{J}(\Gamma),$ (1)
where $||x_{t+S}-G||$ is the distance between the final state $x_{t+S}$ and $G$, and $\mathcal{J}(\Gamma)$ is a cost function for path planning scaled by a relative weight $\lambda$. In mapless urban environments, the robot pose $x_{t}$ and goal $G$ may be highly noisy, causing $\mathcal{J}(\Gamma)$’s importance to vary dynamically across time.
@@ III-B General Preference-Aligned Cost Functions
To properly define our cost function $\mathcal{J}(\Gamma)$, we first need to introduce the notion of an observation function $\Theta:\mathcal{O}\rightarrow\mathcal{T}$ that maps observations $o_{t}$ to a joint embedding space $\mathcal{T}$ that encapsulates a set of relevant factors for navigation in the world. In general, the sufficient set of factors is unknown but can be approximated for each environment. Let $T:\mathcal{X}\rightarrow\mathcal{T}$ be a function that maps a robot pose $x\in\mathcal{X}$ to a feature $\tau\in\mathcal{T}$, where $\tau$ captures the relevant set of factors necessary to reason about $\mathcal{J}(\Gamma)$. The relationship and set of relevant factors may consist of geometric, semantic, and social costs as follows:
@@
where $\sigma$ is a function that combines individual cost terms.
@@ IV Approach
Figure 2: Model architecture and training procedure for the CREStE perceptual encoder $\Theta$. Using a RGB and sparse depth image, $f_{\text{rgbd}}$ extracts image features $z_{\text{rgbd}}$ and performs depth completion. Next, we lift and splat $z_{\text{rgbd}}$ to an unstructured BEV feature map before predicting continuous static panoptic, dynamic panoptic, and elevation features. Finally, we stack the predicted features to construct a structured BEV feature map for our learned reward function. We supervise $\Theta$ using semantic feature maps, completed depth, and BEV map labels generated by our SegmentAnythingv2 [17] and Dinov2 [25]-powered distillation label generator. We define the dimensions for important feature maps in the Feature Dimension Legend on the bottom right.
CREStE is a modular approach with two key components that can be trained end-to-end: 1) A perceptual encoder $\Theta(o_{\text{rgb, t}}$, $o_{\text{depth, t}})$ that takes the robot’s current RGB and sparse depth observation and predicts a completed depth image $y_{\text{depth, t}}$ and structured BEV feature map $y_{\text{bev, t}}$; 2) A reward function $r_{\phi}(y_{\text{bev, t}})$ that takes $y_{\text{bev, t}}$ and outputs a BEV scalar reward map $y_{\text{reward, t}}$. In the remainder of this section, we describe the following: 1) Sec. IV-A - The CREStE model architecture and 2) Sec. IV-B - The CREStE training procedure, where Sec. IV-B1 presents our VFM distillation objective for $\Theta$ and Sec. IV-B2 presents counterfactual IRL and our active reward learning framework for learning $r_{\phi}$.
@@ IV-A1 Perceptual Encoder Model Architecture
Our 25.5M parameter perceptual encoder $\Theta$ draws inspiration from the TerrainNet [21] backbone, which trains a RGB-D encoder $f_{\text{rgbd}}$ to predict a latent feature map $z_{\text{rgbd}}$ using an EfficientNet-B0 [42] encoder and a completed depth map $y_{\text{depth}}$ using a depth completion head $f_{\text{depth}}(z_{\text{rgbd}})$. Like TerrainNet, we train a lift-splat module $f_{\text{splat}}(z_{\text{rgbd}},y_{\text{depth}})$ to lift latent features to 3D and “splat” them to an unstructured BEV feature map $z_{\text{bev, splat}}$. Finally, we pass $z_{\text{bev, splat}}$ to a BEV inpainting backbone $f_{\text{bev}}$ that uses a shared U-Net [32] encoder and separate decoders to predict a structured feature map consisting of separate semantic and elevation layers.
Building on TerrainNet, we make two key architectural modifications. Our first modification, semantic decoder $f_{\text{semantic}}$, promotes learning semantic and geometric-aware features by regressing image features from Dinov2 [25] using latent image features $z_{\text{rgbd}}$. This design is analogous to model distillation [29] and allows our RGB-D encoder $f_{\text{rgbd}}$ to inherit properties from VFMs like robustness to perceptual aliasing and open-set semantic understanding.
Our second modification stems from the observation that Dinov2 features alone lack the entity understanding needed to backproject and inpaint features in heavily occluded urban scenes with noisy depth predictions. Thus, we supplement $f_{\text{bev}}$ with two panoptic map decoders $f_{\text{bev,static}}$ and $f_{\text{bev,dynamic}}$ to ensure that predicted BEV feature maps $y_{\text{bev}}$ are consistent with BEV instance maps from SegmentAnythingv2 (SAM2) [30]. We use Supervised Contrastive Loss [16] to optimize $y_{\text{bev}}$ such that features belonging to the same instance are closer in embedding space than features belonging to different instances. Combined with $f_{\text{semantic}}$, our modifications synergistically unify the strengths of Dinov2 and SAM2 to learn open-set semantic, geometric, and instance-aware representations grounded in the local planning horizon. Altogether, our structured BEV feature map $y_{\text{bev}}$ consists of three layers stacked along the channel dimension: 1) Static panoptic feature map $y_{\text{bev, static}}$, 2) Dynamic panoptic feature map $y_{\text{bev, dynamic}}$, and 3) Elevation map $y_{\text{bev, elev}}$.